OpenStudy (anonymous):

A triangular prism has a base with an area of 45 square inches. The prism also has a length of 6 1/2 inches. Find the volume of this triangular prism.

OpenStudy (anonymous):

I'll give medal to anyone that solves

OpenStudy (amistre64):

multiplication is pretty simple ...

OpenStudy (anonymous):

My brain is pooped i was just helping a guy in calculus which i ain't even taking yet, so yeah, simple things are eluding me

OpenStudy (amistre64):

the dimensions of your area are off, area is a ^2 dimension (im assuming a mistype?)

OpenStudy (anonymous):

The volume V of a prism of base area A and height h is: V = Ah / 3 = 45 * (13 / 2) / 3 = 97.5 in^3.

OpenStudy (anonymous):

thanks Syeda

OpenStudy (anonymous):

No problem

OpenStudy (amistre64):

thats not correct ....

OpenStudy (amistre64):

a prism is not a cone ...

OpenStudy (anonymous):

Oh :o Then yeah Idk.